Community Hospital

Annual Statistics 20XX

INPATIENT DISCHARGES AND DEATHS:

Adults and children

8,234

Newborn

820

INPATIENT DEATHS (included in discharges):

Adults and children

54

Newborn

3

FETAL DEATHS:

Early

4

Intermediate

5

Late

3

INPATIENT AUTOPSIES:

Adults and children

12

Newborn

2

CORONER’S CASES:

Unavailable for autopsy

4

Autopsy by hospital pathologist (included in inpatient autopsies)

1

FETAL DEATH AUTOPSIES:

Intermediate and late

4

  1. What is the adjusted hospital autopsy rate
    1. 29.09%
    2. 2.90%
    3. 29.09%
    4. d.0.29%
  2. What is the fetal autopsy rate?
    1. 0.50%
    2. 5.00%
    3. 50.00%
    4. 0.050%

Please show steps
Answer #1

Adjusted hospital autopsy rate =Total hospital autopsies x 100 / Total number of deaths of hospital patients whose bodies are available for autopsy

Former patient brought in for autopsy = 2

Total number of hospital autopsies= 12(adults&children)+2(newborn)+2(former patient )=16

Total number of deaths of hospital patients whose bodies are available for autopsy= 54+3+2-4(coroner cases)=55

Adjusted hospital autopsy rate = 16 x 100/55=1600/55= 29.09% Option a & c (both options are same)

Fetal autopsy rate = The number of autopsies performed on intermediate and late fetal deaths for a given time period x 100 /the total number of intermediate and late fetal deaths for the same time period

The number of autopsies performed on intermediate and late fetal deaths for a given time period = 4

The total number of intermediate and late fetal deaths for the same time period = 5+3=8

= 4x 100/8

=400/8= 50.00% Option C,
Answer #2

University hospital September 20xx satistics. Inpatiet discharges and deaths: Adults and children 1432 newborn 123... what adjusted hospital autopsy rate?